If a boat goes downstream 72 miles in 3 hours and upstream 60 miles in 6 hours, the rate of the river and the rate of the boat in still water respectively are ____ ?
Answer by Alan3354(69443)

72/3 = 24 mph
60/6 = 10 mph
The boat's speed is the average = (24 + 10)/2 = 17 mph
The current speed is the difference = 7 mph